# What Shape is the Inflationary Bispectrum?

### 17 Mar 2026 v1

@@ Abstract
Observations of the Cosmic Microwave Background are consistent with an inflationary primordial epoch, featuring an almost-exponential background expansion combined with quantum fluctuations around a Bunch Davies vacuum [e.g., 80, 6]. A central prediction of this model is a nearly scale-invariant spectrum of primordial curvature fluctuations, defined by the two-point function $P_{\zeta}(k)\propto k^{n_{s}-4}$ where $n_{s}$ is close to unity [35, 47, 51, 50, 15]. Practically, this implies that the background geometry of the Universe is almost de Sitter, with corrections suppressed by slow-roll parameters. Asserting that the inflationary Lagrangian also (almost) obeys de Sitter isometries greatly restricts the space of inflationary models, for example ensuring that masses and coupling coefficients are time-invariant. Importantly, this assumption implies that all inflationary correlators are approximately scale-invariant, with, for example, the bispectrum, $B_{\zeta}$ scaling as $\alpha^{-6}$ under the dilation $k\to\alpha k$. For this reason, it is common to define a dimensionless shape function $S(x,y)\sim k^{6}B_{\zeta}(kx,ky,k)$ which is a function only of momentum ratios [48, 11, 31].
@@
A particularly interesting example is the primordial exchange of massive particles, which leads to oscillatory signals in the squeezed limit of $S(x,y)$ (with $x\ll y\approx 1$), whose frequency and angular dependence encode the mass and spin of the exchange particle respectively [e.g., 22, 23, 46, 10, 19]; this facilitates particle spectrometry for masses around the inflationary Hubble scale (which could exceed $10^{13}\,\mathrm{GeV}$ [2]). Despite the simple form of the correlator in the squeezed limit, it has become clear that a full analysis of the exchange interactions requires going beyond the squeezed limits, and analyzing the full momentum dependence of the correlator [e.g., 65, 39, 86, 46].
@@
If new physics is lurking in the inflationary three-point function, how can we detect it? To date, most analyses of primordial non-Gaussianity (PNG) focus on the CMB temperature and polarization anisotropies, which currently dominate the inflationary information content [e.g., 73, 5, 41, 79, 49, 27], though large-scale structure is beginning to yield competitive results [25]. Despite its linear nature, the CMB is difficult to analyze, owing to its high dimensionality ($\mathcal{O}(10^{7})$ pixels) and oscillatory transfer functions, which relate the observed anisotropies to primordial curvature [e.g., 43, 44]. Typically, one searches for inflationary signatures by performing targeted searches for specific models of interest, rather than attempting to measure a set of model-independent summary statistics that can then be compared to theoretical models [e.g., 5]. Many of the strongest constraints on inflationary models arise from Komatsu-Spergel-Wandelt estimators [45, 43, 44, 77], which compress the full set of $\mathcal{O}(10^{21})$ bispectrum configurations into a single number. Although quasi-optimal, this procedure is computationally demanding and must be repeated for every primordial bispectrum of interest. Furthermore, it has limited interpretability, and it remains difficult to assess if there is some residual pattern lurking in the data.
@@
Several methods have been proposed to analyze the CMB three-point function in a more generic fashion, including via computation of the binned harmonic-space bispectrum [14, 59, 60] and compression onto a one-dimensional skew-spectrum [52]. Although such approaches have significant utility in constraining late-Universe signatures [e.g., 26], they are difficult to use as a generic primordial physics probe, since one must perform a costly set of integrals to convert from primordial- to CMB-space [e.g., 55]. A powerful alternative is to perform a basis expansion in primordial space, for example via the modal and cmb-best formalisms [31, 29, 78]. By expressing $B_{\zeta}$ as a polynomial series with coefficients computed from the CMB data, these methods can be used to cheaply constrain arbitrary inflationary models, and have recently used to constrain a variety of particle-exchange scenarios [79, 82, 81].
@@
In this Letter, we go one step further by exploiting the (assumed) de Sitter symmetry of inflation. In particular, we directly estimate the physical quantity predicted by inflationary models: the scale-invariant bispectrum shape, $S^{\rm 2D}(x,y)$. This approach, which practically corresponds to reconstructing the underlying shape in a set of logarithmic bins, has a number of practical benefits. Firstly, we can directly visualize the two-dimensional shape and its squeezed limit (which is home to the collider oscillations); in addition to observational interrogation, this allows one to assess the constraining power of different triangle configurations. Secondly, we can trivially compare theory and data for any scale-invariant model of interest (usually in a few milliseconds). Thirdly, since we work in logarithmic space and restrict to scale-invariant configurations, our basis is efficient and quasi-optimal for realistic inflationary models, including those with fast oscillations in the squeezed limit. Finally, we can build analogous scale-dependent estimators, which allow the data to be interrogated in new ways (for example, assessing the information content as a function of $k$). To showcase our approach, we apply it to Planck data, using the resulting measurements to constrain inflationary particle-exchange using more than 20 00020\,000 models of varying mass, spin, and sound-speed, computed using bootstrap techniques [9, 65, 39]. This opens the door to a wide variety of inflationary studies, which are limited only by the speed of computing theoretical models.
@@
positing translation and rotation invariance. Assuming scale-invariance (with $n_{s}=1$), the power spectrum and bispectrum scale as $P_{\zeta}(k)\sim k^{-3}$ and $B_{\zeta}(k,k,k)\sim k^{-6}$ [e.g., 68], which motivates defining the three-dimensional shape function, $S^{\rm 3D}$:
@@
Since the CMB is a linear tracer of the primordial curvature perturbation, we can estimate $s^{\rm 3D}_{n}$ by correlating the temperature (or polarization) anisotropies at three points in space. As shown in previous works [e.g., 60, 77, 32, 43], the minimum-variance estimator is given by
@@
(omitting the normalization factor), where $\tilde{a}_{\ell m}\equiv[C^{-1}a]_{\ell m}$ is an inverse-variance-weighted CMB map. As shown in the Supplemental Material, the estimator can be efficiently evaluated using a combination of spherical harmonic transforms, Monte Carlo summation, and low-dimensional integration [cf., 77, 62]; this is possible since (Estimators) is factorizable in momentum.
@@
Specifically, we search for bispectra induced by the exchange of weakly-mixed boost-breaking spin-zero, spin-one and spin-two particles, across a spectrum of mass parameters, $\mu$ (up to six times the inflationary Hubble scale), and relative sound-speeds, $c_{s}$. [^4] Due to the availability of bootstrap calculations, we limit ourselves to single-exchange interactions in this Letter. As discussed in [e.g., 22, 20], larger spin-zero signals can be generated from double- and triple-exchange diagrams. In contrast to previous CMB studies, we use the full shape function templates computed using the cosmological bootstrap [65] (rather than simplified approximations), extending the calculation to $c_{s}<1$ following [16] and paying close attention to numerical convergence. For each model, we constrain an overall amplitude $f_{\rm NL}$ (set by the couplings in the effective Lagrangian), encoding the size of the signal in the equilateral limit. In full, we evaluate the shape functions in $171$ triangle configurations, across more than $6000$ combinations of mass and spin and four types of interaction. Whilst the theoretical computation is fairly expensive (requiring $\approx 100$ hours in mathematica), the application to data is not: we obtain all of the $f_{\rm NL}$ measurements and errorbars in just $0.6\,\mathrm{s}$.
@@
Figure 3: Constraints on the exchange of massive particles in inflation. We show results for four choices of interaction (induced by the cubic vertices shown in the titles), as a function of mass parameter $\mu\equiv\sqrt{m^{2}_{\sigma}/H^{2}-9/4}$ (or $\sqrt{m^{2}_{\sigma}/H^{2}-(s-1/2)^{2}}$ for spin $s>0$), relative sound-speed $c_{s}=c_{\pi}/c_{\sigma}$, and spin (zero, zero, one, and two respectively). For higher-spin exchange, we restrict to $\mathrm{Im}[\mu]\leq 1/2$ according to the Higuchi bound [36]. The theoretical models are computed numerically using bootstrap techniques (in contrast to most previous works, which used simplified forms, often focussing on the non-oscillatory contribution). The top panel shows the signal-to-noise ratio, whilst the bottom gives the $1\sigma$ error on the $f_{\rm NL}$ amplitude. We report no detection of collider non-Gaussianity with a maximum detection significance of $2.6\sigma$ (for spin-two with $\mu\approx 3,c_{s}\approx 5$), and place strong constraints on the models across a broad range of masses and spins. Given the shape function measurements from Fig. 1, computation of these constraints required just $0.6$ seconds.
@@ Discussion
The above results elucidate an exciting opportunity for inflationary cosmology: using the reconstructed shape functions, we can quickly and accurate place constraints on any scale-invariant inflationary model. In this Letter, we have considered a single class of models, but there are many others worthy of exploration, including inflationary self-interactions [e.g., 18, 17, 24, 73], multiple exchanges [e.g., 46, 87], strongly-mixed sectors [e.g., 86, 67], strongly-coupled sectors [e.g., 8, 66, 40, 34], chemical potentials [e.g., 21, 84, 13], loop diagrams [e.g., 74, 88, 69, 12], and dissipative effects [e.g., 72]. Since we do not require that the input templates are factorizable, we can directly analyze numerically-computed bispectra, allowing connection with recent inflationary developments including the cosmological bootstrap [e.g., 9, 53], Mellin representations [e.g., 69, 70, 76], and flow-based methods [e.g., 67]. Finally, one can analyze families of models simultaneously; this could facilitate a direct analysis of the effective inflationary Lagrangian, accounting for non-linearly realized symmetries.
@@
Finally, the methodology from this Letter could be developed in new directions. A straightforward extension would be to the tensor shape function [cf., 28, 55, 56]. This would greatly simplify analyses of various higher-spin inflationary models and could be achieved by adding polarization factors and exchanging scalar transfer functions for their tensor equivalents. Even more excitingly, one could build a binned estimator for the scale-invariant inflationary trispectrum (building on the modal decompositions of [30]). Since the four-point function is a five-dimensional object, this is highly non-trivial, but could yield important constraints on inflationary scattering processes. We leave these extensions (and many others) to the future.
@@ I.0.1 Three Dimensions
[e.g., 62, 60, 77, 31, 32]. Here, $\tilde{a}\equiv\mathsf{S}^{-1}\,a^{\rm obs}$ for some weighting operator $\mathsf{S}^{-1}$, and $\mathsf{P}$ is the pointing matrix, encoding the response of the observed field $a^{\rm obs}$ to the underlying CMB realization $a$: $a^{\rm obs}=\mathsf{P}a+\text{noise}$. This is naught but a matched filter applied to the $\mathsf{S}^{-1}$-weighted data, with the normalization fixed by requiring that the estimator is unbiased for arbitrary $\mathsf{S}^{-1}$. Setting $\mathsf{S}^{-1}\mathsf{P}$ equal to the inverse covariance matrix of $a^{\rm obs}$, the estimator satisfies its Cramér-Rao bound for $\{s_{n}^{\rm 3D}\}$, achieving minimum-variance. In this limit, the estimator has covariance $\mathcal{F}^{-1,\rm 3D}$, implying that $\mathcal{F}^{\rm 3D}$ is the (mask- and beam-dependent) Fisher matrix.
@@
with permutations over the bin orderings. Practically, computation of $\mathcal{F}^{\rm 3D}$ reduces to a set of harmonic transforms and low-dimensional integrals, following a sum over a small set of Monte Carlo maps $\{\alpha^{(i)}\}$. This form of $\mathcal{F}$ fully accounts for the observational mask, beam, and noise properties, going beyond the simplified diagonal limits used in most CMB estimators [e.g., 78, 29].
@@ II.0.3 Constraints on Standard Templates
where $\mathrm{Cov}_{NN^{\prime}}$ is the covariance of the binned shape function measurements, and $\{S^{{\rm theory},(i)}\}$ are the theoretical models, evaluated at the bin-centers and normalized to unity at $x=y=1$. Here, we consider several options for the covariance: (1) the inverse Fisher matrix, $\mathcal{F}^{-1}$; (2) the sample covariance from the FFP10 simulations; (3) a hybrid approach, combining the correlation structure of $\mathcal{F}^{-1}$ with the amplitude of the empirical covariance (obtained by rotating the FFP10 measurements into a diagonal basis using the Cholesky factorization of $\mathcal{F}$). Whilst option (2) includes any sources of suboptimality (sourced, for example, by our assumption of a diagonal weighting scheme), it is limited by the number of simulations (here $400$ for a $171$-dimensional statistic), thus we primarily adopt option (3) in this study (following a number of previous works [e.g., 58, 55]).

### 26 Mar 2026 v2

@@ Abstract
Observations of the Cosmic Microwave Background are consistent with an inflationary primordial epoch, featuring an almost-exponential background expansion combined with quantum fluctuations around a Bunch Davies vacuum [e.g., 83, 6]. A central prediction of this model is a nearly scale-invariant spectrum of primordial curvature fluctuations, defined by the two-point function $P_{\zeta}(k)\propto k^{n_{s}-4}$ where $n_{s}$ is close to unity [35, 48, 52, 51, 15]. Practically, this implies that the background geometry of the Universe is almost de Sitter, with corrections suppressed by slow-roll parameters. Asserting that the inflationary Lagrangian also (almost) obeys de Sitter isometries greatly restricts the space of inflationary models, for example ensuring that masses and coupling coefficients are time-invariant. Importantly, this assumption implies that all inflationary correlators are approximately scale-invariant, with, for example, the bispectrum, $B_{\zeta}$ scaling as $\alpha^{-6}$ under the dilation $k\to\alpha k$. For this reason, it is common to define a dimensionless shape function $S(x,y)\sim k^{6}(xy)^{2}B_{\zeta}(kx,ky,k)$ which is a function only of momentum ratios [49, 11, 31].
@@
A particularly interesting example is the primordial exchange of massive particles, which leads to oscillatory signals in the squeezed limit of $S(x,y)$ (with $x\ll y\approx 1$), whose frequency and angular dependence encode the mass and spin of the exchange particle respectively [e.g., 22, 23, 47, 10, 19]; this facilitates particle spectrometry for masses around the inflationary Hubble scale (which could exceed $10^{13}\,\mathrm{GeV}$ [2]). Despite the simple form of the correlator in the squeezed limit, it has become clear that a full analysis of the exchange interactions requires going beyond the squeezed limits, and analyzing the full momentum dependence of the correlator [e.g., 66, 39, 90, 47].
@@
If new physics is lurking in the inflationary three-point function, how can we detect it? To date, most analyses of primordial non-Gaussianity (PNG) focus on the CMB temperature and polarization anisotropies, which currently dominate the inflationary information content [e.g., 75, 5, 41, 81, 50, 27], though large-scale structure is beginning to yield competitive results [25]. Despite its linear nature, the CMB is difficult to analyze, owing to its high dimensionality ($\mathcal{O}(10^{7})$ pixels) and oscillatory transfer functions, which relate the observed anisotropies to primordial curvature [e.g., 43, 44]. Typically, one searches for inflationary signatures by performing targeted searches for specific models of interest, rather than attempting to measure a set of model-independent summary statistics that can then be compared to theoretical models [e.g., 5]. Many of the strongest constraints on inflationary models arise from Komatsu-Spergel-Wandelt estimators [45, 43, 44, 79], which compress the full set of $\mathcal{O}(10^{21})$ bispectrum configurations into a single number. Although quasi-optimal, this procedure is computationally demanding and must be repeated for every primordial bispectrum of interest. Furthermore, it has limited interpretability, and it remains difficult to assess if there is some residual pattern lurking in the data.
@@
Several methods have been proposed to analyze the CMB three-point function in a more generic fashion, including via computation of the binned harmonic-space bispectrum [14, 60, 61] and compression onto a one-dimensional skew-spectrum [53]. Although such approaches have significant utility in constraining late-Universe signatures [e.g., 26], they are difficult to use as a generic primordial physics probe, since one must perform a costly set of integrals to convert from primordial- to CMB-space [e.g., 57]. A powerful alternative is to perform a basis expansion in primordial space, for example via the modal and cmb-best formalisms [31, 29, 80]. By expressing $B_{\zeta}$ as a polynomial series with coefficients computed from the CMB data, these methods can be used to cheaply constrain arbitrary inflationary models, and have recently used to constrain a variety of particle-exchange scenarios [81, 85, 84, 46, 74].
@@
In this Letter, we go one step further by exploiting the (assumed) dilatation symmetry of inflation. In particular, we directly estimate the physical quantity predicted by inflationary models: the scale-invariant bispectrum shape, $S^{\rm 2D}(x,y)$. This approach, which practically corresponds to reconstructing the underlying shape in a set of logarithmic bins, has a number of practical benefits. Firstly, we can directly visualize the two-dimensional shape and its squeezed limit (which is home to the collider oscillations); in addition to observational interrogation, this allows one to assess the constraining power of different triangle configurations. Secondly, we can trivially compare theory and data for any scale-invariant model of interest (usually in a few milliseconds). Thirdly, since we work in logarithmic space and restrict to scale-invariant configurations, our basis is efficient and quasi-optimal for realistic inflationary models, including those with fast oscillations in the squeezed limit. Finally, we can build analogous scale-dependent estimators, which allow the data to be interrogated in new ways (for example, assessing the information content as a function of $k$). To showcase our approach, we apply it to Planck data, using the resulting measurements to constrain inflationary particle-exchange using more than 20 00020\,000 models of varying mass, spin, and sound-speed, computed using bootstrap techniques [9, 66, 39]. This opens the door to a wide variety of inflationary studies, which are limited only by the speed of computing theoretical models.
@@
positing translation and rotation invariance. Assuming scale-invariance (with $n_{s}=1$), the power spectrum and bispectrum scale as $P_{\zeta}(k)\sim k^{-3}$ and $B_{\zeta}(k,k,k)\sim k^{-6}$ [e.g., 69], which motivates defining the three-dimensional shape function, $S^{\rm 3D}$:
@@
Since the CMB is a linear tracer of the primordial curvature perturbation, we can estimate $s^{\rm 3D}_{n}$ by correlating the temperature (or polarization) anisotropies at three points in space. As shown in previous works [e.g., 61, 79, 32, 43], the minimum-variance estimator is given by
@@
(omitting the normalization factor), where $\tilde{a}_{\ell m}\equiv[C^{-1}a]_{\ell m}$ is an inverse-variance-weighted CMB map. As shown in the Supplemental Material, the estimator can be efficiently evaluated using a combination of spherical harmonic transforms, Monte Carlo summation, and low-dimensional integration [cf., 79, 63]; this is possible since (Estimators) is factorizable in momentum.
@@
Specifically, we search for bispectra induced by the exchange of weakly-mixed boost-breaking spin-zero, spin-one and spin-two particles, across a spectrum of mass parameters, $\mu$ (up to six times the inflationary Hubble scale), and relative sound-speeds, $c_{s}$. [^4] Due to the availability of bootstrap calculations, we limit ourselves to single-exchange interactions in this Letter. As discussed in [e.g., 22, 20, 46], larger spin-zero signals can be generated from double- and triple-exchange diagrams. In contrast to previous CMB studies, we use the full shape function templates computed using the cosmological bootstrap [66] (rather than simplified approximations), extending the calculation to $c_{s}<1$ following [16] and paying close attention to numerical convergence. For each model, we constrain an overall amplitude $f_{\rm NL}$ (set by the couplings in the effective Lagrangian), encoding the size of the signal in the equilateral limit. In full, we evaluate the shape functions in $171$ triangle configurations, across more than $6000$ combinations of mass and spin and four types of interaction. Whilst the theoretical computation is fairly expensive (requiring $\approx 100$ hours in mathematica), the application to data is not: we obtain all of the $f_{\rm NL}$ measurements and errorbars in just $0.6\,\mathrm{s}$.
@@
Figure 3: Constraints on the exchange of massive particles in inflation. We show results for four choices of interaction (induced by the cubic vertices shown in the titles; see the Supplemental Material for the underlying Lagrangian), as a function of mass parameter $\mu\equiv\sqrt{m^{2}_{\sigma}/H^{2}-9/4}$ (or $\sqrt{m^{2}_{\sigma}/H^{2}-(s-1/2)^{2}}$ for spin $s>0$), relative sound-speed $c_{s}=c_{\pi}/c_{\sigma}$, and spin (zero, zero, one, and two respectively). For higher-spin exchange, we restrict to $\mathrm{Im}[\mu]\leq 1/2$ according to the Higuchi bound [36]. The theoretical models are computed numerically using bootstrap techniques (in contrast to most previous works, which used simplified forms, often focussing on the non-oscillatory contribution). The top panel shows the signal-to-noise ratio, whilst the bottom gives the $1\sigma$ error on the $f_{\rm NL}$ amplitude. We report no detection of collider non-Gaussianity with a maximum detection significance of $2.6\sigma$ (for spin-two with $\mu\approx 3,c_{s}\approx 5$), and place strong constraints on the models across a broad range of masses and spins. Given the shape function measurements from Fig. 1, computation of these constraints required just $0.6$ seconds.
@@ Discussion
The above results elucidate an exciting opportunity for inflationary cosmology: using the reconstructed shape functions, we can quickly and accurate place constraints on any scale-invariant inflationary model. In this Letter, we have considered a single class of models, but there are many others worthy of exploration, including inflationary self-interactions [e.g., 18, 17, 24, 75], multiple exchanges [e.g., 47, 91], strongly-mixed sectors [e.g., 90, 68], strongly-coupled sectors [e.g., 8, 67, 40, 34], chemical potentials [e.g., 21, 88, 13, 82, 86], loop diagrams [e.g., 76, 92, 70, 12], and dissipative effects [e.g., 73]. Since we do not require that the input templates are factorizable, we can directly analyze numerically-computed bispectra, allowing connection with recent inflationary developments including the cosmological bootstrap [e.g., 9, 54], Mellin representations [e.g., 70, 71, 78], and flow-based methods [e.g., 68]. Finally, one can analyze families of models simultaneously; this could facilitate a direct analysis of the effective inflationary Lagrangian, accounting for non-linearly realized symmetries.
@@
Finally, the methodology from this Letter could be developed in new directions. A straightforward extension would be to the tensor shape function [cf., 28, 57, 56]. This would greatly simplify analyses of various higher-spin inflationary models and could be achieved by adding polarization factors and exchanging scalar transfer functions for their tensor equivalents. Even more excitingly, one could build a binned estimator for the scale-invariant inflationary trispectrum (building on the modal decompositions of [30]). Since the four-point function is a five-dimensional object, this is highly non-trivial, but could yield important constraints on inflationary scattering processes. We leave these extensions (and many others) to the future.
@@ I.0.1 Three Dimensions
[e.g., 63, 61, 79, 31, 32]. Here, $\tilde{a}\equiv\mathsf{S}^{-1}\,a^{\rm obs}$ for some weighting operator $\mathsf{S}^{-1}$, and $\mathsf{P}$ is the pointing matrix, encoding the response of the observed field $a^{\rm obs}$ to the underlying CMB realization $a$: $a^{\rm obs}=\mathsf{P}a+\text{noise}$. This is naught but a matched filter applied to the $\mathsf{S}^{-1}$-weighted data, with the normalization fixed by requiring that the estimator is unbiased for arbitrary $\mathsf{S}^{-1}$. Setting $\mathsf{S}^{-1}\mathsf{P}$ equal to the inverse covariance matrix of $a^{\rm obs}$, the estimator satisfies its Cramér-Rao bound for $\{s_{n}^{\rm 3D}\}$, achieving minimum-variance. In this limit, the estimator has covariance $\mathcal{F}^{-1,\rm 3D}$, implying that $\mathcal{F}^{\rm 3D}$ is the (mask- and beam-dependent) Fisher matrix.
@@
with permutations over the bin orderings. Practically, computation of $\mathcal{F}^{\rm 3D}$ reduces to a set of harmonic transforms and low-dimensional integrals, following a sum over a small set of Monte Carlo maps $\{\alpha^{(i)}\}$. This form of $\mathcal{F}$ fully accounts for the observational mask, beam, and noise properties, going beyond the simplified diagonal limits used in most CMB estimators [e.g., 80, 29].
@@ II.0.3 Constraints on Standard Templates
where $\mathrm{Cov}_{NN^{\prime}}$ is the covariance of the binned shape function measurements, and $\{S^{{\rm theory},(i)}\}$ are the theoretical models, evaluated at the bin-centers and normalized to unity at $x=y=1$. Here, we consider several options for the covariance: (1) the inverse Fisher matrix, $\mathcal{F}^{-1}$; (2) the sample covariance from the FFP10 simulations; (3) a hybrid approach, combining the correlation structure of $\mathcal{F}^{-1}$ with the amplitude of the empirical covariance (obtained by rotating the FFP10 measurements into a diagonal basis using the Cholesky factorization of $\mathcal{F}$). Whilst option (2) includes any sources of suboptimality (sourced, for example, by our assumption of a diagonal weighting scheme), it is limited by the number of simulations (here $400$ for a $171$-dimensional statistic), thus we primarily adopt option (3) in this study (following a number of previous works [e.g., 59, 57]).
@@ II.0.4 Scale Dependence
Multi-Field Lagrangian
Here, we specify the inflationary Lagrangian used to compute the exchange bispectra shown in Fig. 3. In terms of the Goldstone mode of broken time translations, $\pi$, the basic Lagrangian is given by
$a^{-3}\mathcal{L}$ $\supset$ $\frac{M_{\rm Pl}^{2}|\dot{H}|}{c_{s}^{2}}\left(\dot{\pi}^{2}-c_{s}^{2}\frac{(\partial_{i}\pi)^{2}}{a^{2}}\right)+a^{-3}\mathcal{L}_{\pi}^{(\rm int)},$ (S18)
where $H$ is the Hubble parameter, $c_{s}$ is the sound-speed, and $\mathcal{L}_{\pi}^{(\rm int)}$ represents higher-order self-interactions that we ignore in this study. Introducing a scalar field, $\sigma$ (optionally with spin, but setting $c_{\sigma}=1$ without loss of generality), we induce new terms in the Lagrangian:
$a^{-3}\mathcal{L}$ $\supset$ $a^{-3}\mathcal{L}_{\sigma}^{(2)}+\begin{cases}\rho_{0}\dot{\pi}\sigma+\lambda_{0}\dot{\pi}^{2}\sigma\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+a^{-2}\lambda_{0}^{\prime}(\partial_{i}\pi)^{2}\sigma&\text{spin-0}\\ a^{-2}\left(\rho_{1}\partial_{i}\pi\sigma^{i}+\lambda_{1}\dot{\pi}\partial_{i}\pi\sigma^{i}\right)&\text{spin-1}\\ a^{-4}\left(\rho_{2}\partial_{ij}\pi\hat{\sigma}^{ij}+\lambda_{2}\dot{\pi}\partial_{ij}\pi\hat{\sigma}^{ij}\right)&\text{spin-2},\end{cases}$ (S19)
where $\mathcal{L}_{\sigma}^{(2)}$ is the free-field quadratic Lagrangian. We have dropped terms that contribute only to double- and triple-exchange diagrams (e.g., $\dot{\pi}\sigma^{2}$), though these can source large interaction signatures [e.g., 46]. In the main text, we constrain the $f_{\rm NL}$ amplitudes corresponding to each of the four cubic interactions above, with $f_{\rm NL}\propto\rho_{n}\lambda_{n}$.